OpenStudy (nickymail48):

how to solve quadritic formala then simplify? x^2+8x+14=0

OpenStudy (johnweldon1993):

So you need to figure out: 1) what 2 numbers multiply to make 14 AND 2) do those 2 numbers ALSO add to make 8?

OpenStudy (nickymail48):

so i write as quadritic formula right?

OpenStudy (triciaal):

do you already know the formula?

OpenStudy (triciaal):

@johnweldon1993 not solved by factoring hence required to use the formula

OpenStudy (nickymail48):

yes i do but then its wrong cause they wan me to simplify

OpenStudy (johnweldon1993):

You can via grouping as I did, however, since you (the poster) brought up the quadratic formula, absolutely we can do that \[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (johnweldon1993):

What is 'a' 'b' and 'c' in this equation?

OpenStudy (nickymail48):

1,8,14

OpenStudy (johnweldon1993):

Great, so we have \[\large \frac{-8 \pm \sqrt{8^2 - 4(1)(14)}}{2(1)}\] The painful part is the simplification \[\large \frac{-8 \pm \sqrt{64 - 56}}{2}\] what does THAT simplify to?

OpenStudy (nickymail48):

-8+8/2

OpenStudy (johnweldon1993):

Not quite, you lost the square root \[\large \frac{-8 \pm \sqrt{8}}{2}\] now, we can simplify that further right? what if I wrote this like \[\large \frac{-8 \pm \sqrt{4\times 2}}{2}\] what could we do?

OpenStudy (nickymail48):

divide the 8 by 2?

OpenStudy (johnweldon1993):

Not quite, was mainly talking about the square root part Hint* what is the square root of 4?

OpenStudy (nickymail48):

2

OpenStudy (johnweldon1993):

Right so we can change \[\large \frac{-8 \pm \sqrt{4\times 2}}{2}\] to \[\large \frac{-8 \pm 2\sqrt{ 2}}{2}\] NOW we can divide everything by 2 \[\large \frac{\cancel{-8}\rightarrow -4 \pm \cancel{2}\sqrt{2}}{\cancel{2}}\] written more nicely \[\large -4 \pm \sqrt{2}\]

OpenStudy (johnweldon1993):

Make sense?

OpenStudy (nickymail48):

yes but could i just divide the numbers first before solving?

OpenStudy (johnweldon1993):

What do you mean? Like before I simplified the square root?

OpenStudy (nickymail48):

like 8 could be divided to 4 before writing out the problem

OpenStudy (johnweldon1993):

You COULD have, but the 2 would still be in the denominator for the 2nd part of the numerator...see what I mean \[\large \frac{-8 \pm \sqrt{8}}{2}\] \[\large -4 \pm \frac{\sqrt{8}}{2}\]

OpenStudy (johnweldon1993):

That's why I wanted to simplify that square root first....making it 4 times 2...and then bringing the 2 out in front made it so THOSE 2's would cancel and we wouldnt need to deal with that annoying fraction

OpenStudy (nickymail48):

yah so that wont work and isthe anser plus or minus?

OpenStudy (johnweldon1993):

Not quite, Whenever you see a solution with a +/- sign, it means you have 2 equations Here we have \[\large -4 \pm \sqrt{2}\] whih will break into \[\large -4 + \sqrt{2}\] and \[\large -4 - \sqrt{2}\]

OpenStudy (nickymail48):

ok thanks i really got to practice its on a final

OpenStudy (johnweldon1993):

No problem, yeah the quadratic formula can help out a lot in these cases if you're not familiar with grouping or completing the square etc...

OpenStudy (nickymail48):

oh yah i really struggling with factoring trinomials etc